Clark Heter is an industrial engineer at Lyons Products. He would like to determine whether there are more units produced on the night shift than on the day shift. A sample of 60 day-shift workers showed that the mean number of units produced was 334, with a population standard deviation of 23. A sample of 68 night-shift workers showed that the mean number of units produced was 341, with a population standard deviation of 28 units.

At the .10 significance level, is the number of units produced on the night shift larger?
1. This is a (Click to select)twoone-tailed test.
2.

The decision rule is to reject H0: μdμn if z < . (Negative amount should be indicated by a minus sign. Round your answer to 2 decimal places.)

3.

The test statistic is z = . (Negative amount should be indicated by a minus sign. Round your answer to 2 decimal places.)

4. What is your decision regarding H0?
  
(Click to select)Reject.Do not reject.
Answer #1

1)

One tailed test

2)

Rejection Region
This is left tailed test, for α = 0.1
Critical value of z is -1.28
Hence reject H0 if z < -1.28

3)


Pooled Variance
sp = sqrt(s1^2/n1 + s2^2/n2)
sp = sqrt(529/60 + 784/68)
sp = 4.5107

Test statistic,
z = (x1bar - x2bar)/sp
z = (334 - 341)/4.5107
z = -1.55

4)

Reject H0